# Non-Markovian Reduced Models to Unravel Transitions in Non-equilibrium Systems

### v1

@@ VI.1 Spatially extended system with double-fold bifurcation
This system serves as a basic example of a spatially extended system exhibiting a double-fold bifurcation (also known as an S-shaped bifurcation). In this bifurcation, two equilibrium points collide and vanish, leading to tipping phenomena within a hysteresis loop. Similar elliptic problems to Eq. (LABEL:Eq_Sbif) with these properties emerge in various applications such as gas combustion and chemical kinetics [61, 62], plasma physics and Grad-Shafranov equilibria in Tokamaks [130, 63, 64], and gravitational equilibrium of polytropic stars [130, 131, 132]. Beyond polynomial nonlinearities, other spatially extended systems can also exhibit S-shaped bifurcation diagrams. Examples include: Earth’s energy balance models [66, 67, 133], and oceanic models based on hydrostatic primitive or Boussinesq equations employed in the study of the Atlantic meridional overturning circulation (AMOC) [68, 134, 69].
@@
For equation (LABEL:Eq_Sbif), the bifurcation parameter is $\lambda$. For $L=2$, it is known that the solution curve $\lambda\mapsto u_{\lambda}$ exhibits an S-shaped bifurcation when $\epsilon<\epsilon^{*}\approx 0.3103$ [135, Theorem 3.2]. This critical constant $\epsilon^{\ast}$ is derived from a general analytic formula applicable to a broader class of equations where $u^{3}$ is replaced by $u^{p}$ with a fractional exponent $p>2$ [135]. More precisely, for $\epsilon$ satisfying
@@ VI.2 System’s sensitivity to jumps, and jump-driven dynamics
and is solved below using a Chebyshev spectral method [136, Chapters 6 & 7] for the numerical experiments. Hereafter we denote by $(\beta_{n},\bm{e}_{n})$ the eigenpairs solving Eq. (LABEL:Eq_eigenpb). Note that the eigenvalues are real and simple for this problem.
Thus, the jump-driven SPDE that will be at the core of our reduction study is given by:
@@
where $\xi_{n}$ is a uniformly distributed random variable taking values in $[0,1]$ and $\mathds{1}_{\{\xi\leq f_{r}\}}=1$ if and only if $0\leq\xi\leq f_{r}$. We refer to [137, Sec. 3.4] for existence results in Hilbert spaces of mild solutions to SPDEs driven by jump processes, that apply to Eq. (LABEL:Eq_SPDEjump).
@@
This configuration ensures that if a perturbation is applied for a duration of $\Delta t$ and then removed for several subsequent $\Delta t$ intervals, the system will naturally return to its closest stable steady state.
@@ VI.3 Non-Markovian optimal reduced model
We describe below how our Backward-Forward framework introduced in this work to derive stochastic parameterizations of SPDEs driven by white noise can be readily adapted to SPDEs driven by jump processes. Equation (LABEL:Eq_SPDEjump), driven by the jump process defined in Eq. (91), serves as a concrete example. Our objective is to develop a reduced model for the first mode’s amplitude in Eq. (LABEL:Eq_SPDEjump). In other words, we aim to create a simplified, one-dimensional stochastic model that can accurately capture the complex dynamics discussed earlier, including the stochastic transitions.
@@
$\Phi_{n}(\tau_{n},$ $\;\bm{\beta},X,t;\omega)$ (96)
$=e^{\beta_{n}\tau}Y+J^{n}_{\tau}(\bm{\beta},t;\omega)+D^{n}_{11}(\tau,\bm{\beta})B_{11}^{n}X^{2}$
$\;\;+E_{111}^{n}(\tau,\bm{\beta})C_{111}^{n}X^{3},\;n\in\{3,5\},$
@@
$J^{n}_{\tau}(\bm{\beta},t;\omega)=\sigma\lambda\int_{t-\tau}^{t}e^{\beta_{n}(t-s)}\zeta_{s}f(s)\,\mathrm{d}s.$ (97)
@@
for $n=3$ and $n=5$; cf. Eq. (47).
@@
$\mathcal{P}(t)\approx J^{3}_{\tau^{*}_{3}}(\bm{\beta},t;\omega)\bm{e}_{3}+J^{5}_{\tau^{*}_{5}}(\bm{\beta},t;\omega)\bm{e}_{5},$ (100)
@@
To make this point, we introduce a few more notations. We denote by $a_{\lambda}^{\epsilon}(x)$ the spatial coefficient, $\lambda-3\lambda\epsilon{U^{\ast}_{\lambda}}(x)$, involved in the quadratic term of Eq. (LABEL:Eq_reduced_S_shaped) (as inherited from Eq. (94)). To simplify the notations we denote also by $J_{3}^{\ast}$ (resp. $J_{5}^{\ast}$) the term $J^{3}_{\tau^{*}_{3}}(\bm{\beta},t;\omega)$ (resp. $J^{5}_{\tau_{5}^{*}}(\bm{\beta},t;\omega)$), and introduce the bilinear operator $G_{2}(u,v)(x)=u(x)v(x),$ for $x$ in $(0,L)$ and any functions $u,v$ in $L^{2}(0,L)$. We have then, using the approximation formula (100), that the quadratic term in Eq. (LABEL:Eq_reduced_S_shaped) expands as follows
@@
with a similar expression obtained from the cubic terms in Eq. (LABEL:Eq_reduced_S_shaped). Already from (LABEL:Expansion_quadterm) one observes the production of stochastic constant terms driven by $(J_{3}^{\ast}(t))^{2}$, $(J_{5}^{\ast}(t))^{2}$ and $J_{3}^{\ast}(t)J_{5}^{\ast}(t)$ as well as linear terms driven by $J_{3}^{\ast}(t)$ and $J_{5}^{\ast}(t)$ contributing to the dynamics of the surrogate first mode’s amplitude $y$. These stochastic terms turn out to be essential for a good emulation of the first mode’s amplitude [^2] Such Ornstein-Uhlenbeck processes are known to play a key role in rigorous estimates; see e.g. [49], for SPDEs driven by white noise.. This point is illustrated by the following experiment.
@@ VI.4 Training data, parameterization defects, and reduced model’s skills
The training data for the optimal reduced model is comprised of $N=40,000$ SPDE solution profiles. These profiles are generated over a training interval from 00 to $T$, with $T$ = 400, for a single, specific noise path. The corresponding normalized parameterization defects (quantified by $Q_{n}$) are visualized in Figure 13.
@@
Figure 14 shows the optimal parameterization’s ability to captures first mode’s complex behavior. This figure showcases the effectiveness of the optimal reduced model (Eq. (LABEL:Eq_reduced_S_shaped) with $\mathcal{P}(t)$ from Eq. (100)) in capturing the bimodal nature of the first mode’s amplitude. The model’s performance is evaluated using a vast set of unseen test paths. When compared to the actual SPDE solution’s amplitude for the first mode (across these test paths), the optimal reduced model closely follows the various transitions between metastable states, crossing as needed the potential barriers. However, it is worth noting that this reduced model tends to slightly overestimate the intensity of these larger excursions (as shown in Figure 15).
@@
VII Discussion and Perspectives
@@
Thus, this work demonstrates that Optimal Parameterizing Manifolds (OPMs) constructed from Backward-Forward (BF) systems and optimized using a single solution path can capture noise-induced transitions across a broad range of out-of-sample, test paths. This approach is effective for both Gaussian and Lévy-type noise.
@@
The formulas derived from the BF systems in Section III enable practical approximation of the theoretical optimal parameterization, particularly its non-Markovian random coefficients. These coefficients become crucial, especially when noise affects unresolved modes. They essentially capture how noise propagates to the resolved modes through nonlinear interactions.
@@
These more general stochastic parameterizations have the potential to significantly enhance the parameterization capabilities of $\Phi$ (Eq. (33)). This is because they capture higher-order interaction terms with respect to the noise amplitude, as shown in Eq. (30). Resolving these interactions is potentially important for improving the prediction of rare event statistics by non-Markovian reduced systems, especially when the noise intensity is low. This opens doors to potential applications using large deviation or rare event algorithms (e.g., [140, 141, 142, 53, 128, 54, 55, 21]). Additionally, iterating the BF system (Eq. (25)) beyond $\ell=2$ can also introduce higher-order terms involving the low-mode amplitude $X$.
The inclusion of higher-term in $X$ is non-unique though. For instance, inspired by [30, Sec. 7.2.1], stochastic parameterizations accounting for higher-order terms in $X$ are produced by solving analytically the following BF system:
@@
for which the two first equations for $p^{(1)}$ and $p^{(2)}$ are simultaneously integrated backward over $[t-\tau,t]$, followed by the forward integration of the $q_{n}$-equation over the same interval. Compared to Eq. (LABEL:Eq_BF_SPDEcubic), this BF system has two-layer in the low-mode variable $p$ resulting into parameterizations of order 4 in $X$; see [143, Theorem 2] for an example. These higher-order terms come from the nonlinear self-interactions between the low modes as brought by the the term $\Pi_{c}G_{2}\big{(}p^{(1)}\big{)}$. Accounting for additional interactions either through (LABEL:Eq_h2) or the iterative procedure of Section II.3 can help improve the parameterization skills (see also [49, 52]), and should be considered if accuracy is key to resolve.

### v2

@@ VI.1 Spatially extended system with double-fold bifurcation
This system serves as a basic example of a spatially extended system exhibiting a double-fold bifurcation (also known as an S-shaped bifurcation). In this bifurcation, two equilibrium points collide and vanish, leading to tipping phenomena within a hysteresis loop. Similar elliptic problems to Eq. (LABEL:Eq_Sbif) with these properties emerge in various applications such as gas combustion and chemical kinetics [58, 59], plasma physics and Grad-Shafranov equilibria in Tokamaks [137, 60, 61], and gravitational equilibrium of polytropic stars [137, 138, 139]. Beyond polynomial nonlinearities, other spatially extended systems are known to exhibit S-shaped bifurcation diagrams. Examples include: Earth’s energy balance models [63, 64, 140], and oceanic models based on hydrostatic primitive or Boussinesq equations employed in the study of the Atlantic meridional overturning circulation (AMOC) [65, 141, 66]. We also note that the OPM approach has already been successfully applied to predict tipping phenomena in a prototype Stommel–Cessi model of the AMOC driven by Gaussian noise [28].
@@
For equation (LABEL:Eq_Sbif), the bifurcation parameter is $\lambda$. For $L=2$, it is known that the solution curve $\lambda\mapsto u_{\lambda}$ exhibits an S-shaped bifurcation when $\epsilon<\epsilon^{*}\approx 0.3103$ [142, Theorem 3.2]. This critical constant $\epsilon^{\ast}$ is derived from a general analytic formula applicable to a broader class of equations where $u^{3}$ is replaced by $u^{p}$ with a fractional exponent $p>2$ [142]. More precisely, for $\epsilon$ satisfying
@@ VI.2 System’s sensitivity to jumps, and jump-driven dynamics
and is solved below using a Chebyshev spectral method [143, Chapters 6 & 7] for the numerical experiments. Hereafter we denote by $(\beta_{n},\bm{e}_{n})$ the eigenpairs solving Eq. (LABEL:Eq_eigenpb). Note that the eigenvalues are real and simple for this problem.
Thus, the jump-driven SPDE at the core of our reduction study is given by:
@@
where $\xi_{n}$ is a uniformly distributed random variable taking values in $[0,1]$ and $\mathds{1}_{\{\xi\leq f_{r}\}}=1$ if and only if $0\leq\xi\leq f_{r}$. We refer to [83, Sec. 3.4] for existence results in Hilbert spaces of mild solutions to SPDEs driven by jump processes, that apply to Eq. (LABEL:Eq_SPDEjump). This type of jump process is also known as a dichotomous Markov noise [144], or a two-state Markov jump process in certain fields [76]. It is encountered in many applications [145].
@@
This configuration ensures that if a perturbation is applied for a duration of $\Delta t$ and then removed for several subsequent $\Delta t$ intervals, the system naturally return to its closest stable steady state.
@@ VI.3 Non-Markovian optimal reduced model
We describe below how our Backward-Forward framework introduced in this work to derive stochastic parameterizations of SPDEs driven by white noise can be readily adapted to SPDEs driven by jump processes. Equation (LABEL:Eq_SPDEjump), driven by the jump process defined in Eq. (94), serves as a concrete example. Our objective is to develop a reduced model for the first mode’s amplitude in Eq. (LABEL:Eq_SPDEjump). In other words, we aim to create a simplified, one-dimensional stochastic model that can accurately reproduce the complex dynamics discussed in Section VI.2, including the stochastic transitions.
@@
$\Phi_{n}(\tau_{n},$ $X,t;\omega)$ (99)
$=e^{\beta_{n}\tau}Y+J^{n}_{\tau}(t;\omega)+D^{n}_{11}(\tau)B_{11}^{n}X^{2}$
$\;\;+E_{111}^{n}(\tau)C_{111}^{n}X^{3},\;n\in\{3,5\},$
@@
$J^{n}_{\tau}(t;\omega)=\sigma\lambda\int_{t-\tau}^{t}e^{\beta_{n}(t-s)}\zeta_{s}f(s)\,\mathrm{d}s.$ (100)
@@
for $n=3$ and $n=5$; cf. Eq. (49). Note that the parameterization Eq. (99) takes the same structural form as Eq. (82) in the case of white noise forcing (see Remark III.1), with the bounded jump noise replacing the latter here. This is due to the underlying backward-forward systems at the core of the derivation of these formulas where only the nature of the forcing changes; see also Sec. VIII below.
@@
$\mathcal{P}(t)\approx J^{3}_{\tau^{*}_{3}}(t;\omega)\bm{e}_{3}+J^{5}_{\tau^{*}_{5}}(t;\omega)\bm{e}_{5},$ (103)
@@
To make this point, we introduce a few more notations. We denote by $a_{\lambda}^{\epsilon}(x)$ the spatial coefficient, $\lambda-3\lambda\epsilon{U^{\ast}_{\lambda}}(x)$, involved in the quadratic term of Eq. (LABEL:Eq_reduced_S_shaped) (as inherited from Eq. (97)). To simplify the notations we denote also by $J_{3}^{\ast}$ (resp. $J_{5}^{\ast}$) the term $J^{3}_{\tau^{*}_{3}}(t;\omega)$ (resp. $J^{5}_{\tau_{5}^{*}}(t;\omega)$), and introduce the bilinear operator $G_{2}(u,v)(x)=u(x)v(x),$ for $x$ in $(0,L)$ and any functions $u,v$ in $L^{2}(0,L)$. We have then, using the approximation formula (103), that the quadratic term in Eq. (LABEL:Eq_reduced_S_shaped) expands as follows
@@
with a similar expression obtained from the cubic terms in Eq. (LABEL:Eq_reduced_S_shaped). Already from (LABEL:Expansion_quadterm) one observes the production of stochastic constant terms driven by $(J_{3}^{\ast}(t))^{2}$, $(J_{5}^{\ast}(t))^{2}$ and $J_{3}^{\ast}(t)J_{5}^{\ast}(t)$ as well as linear terms driven by $J_{3}^{\ast}(t)$ and $J_{5}^{\ast}(t)$ contributing to the dynamics of the surrogate first mode’s amplitude $y$. These stochastic terms turn out to be essential for a good emulation of the first mode’s amplitude. This point is illustrated by the following experiment.
@@ VI.4 Training data, parameterization defects, and reduced model’s skills
The training data for the optimal reduced model is comprised of $N=40,000$ snapshots. These sACE snapshots are generated over a training interval from 00 to $T$, with $T$ = 400, for a single, noise path driving the sACE. The corresponding normalized parameterization defects (quantified by $Q_{n}$) are visualized in Figure 13.
@@
Figure 14 shows the optimal parameterization’s ability to reproduce first mode’s complex behavior. This figure showcases the effectiveness of the optimal reduced model (Eq. (LABEL:Eq_reduced_S_shaped) with $\mathcal{P}(t)$ from Eq. (103)) in capturing the bimodal nature of the first mode’s amplitude. The model’s performance is evaluated using a vast set of unseen test paths. When compared to the actual SPDE solution’s amplitude for the first mode (across these test paths), the optimal reduced model closely follows the various transitions between metastable states, crossing as needed the potential barriers. However, it is worth noting that this reduced model tends to slightly overestimate the intensity of these larger excursions (as shown in Figure 15).
@@
VII Towards the Reduction of More General Lévy-driven Dynamics
We prolong the discussion and results of Section VI towards more general Lévy noise. For the sake of simplicity we focus on finite-dimensional SDEs driven by Lévy processes.
There is a vast literature on Lévy processes which roughly speaking are processes given as the sum of a Brownian motion and a jump process [147, 148, 149, 150, 151]. Unlike the case of diffusion processes (SDEs driven by Brownian motions), the representation of Kolmogorov operators is non-unique in the case of Lévy-driven SDEs, and may take the form of operators involving fractional Laplacians or singular integrals among other representations [152]. We adopt here the definition commonly used in probability theory [153, Theorem 3.5.14] which presents the interest of being particularly intuitive as recalled below. We consider SDEs of the form
$\,\mathrm{d}X_{t}={\bm{F}}(X_{t})\,\mathrm{d}t+{\bm{\Sigma}}(X_{t})\,\mathrm{d}{\bm{W}_{t}}+\,\mathrm{d}\bm{L}_{t},$ (105)
where ${\bm{W}_{t}}$ is a Brownian motion in $\mathbb{R}^{d}$, and $\bm{L}_{t}$ is a Lévy process on $\mathbb{R}^{d}$ independent from ${\bm{W}_{t}}$. Here, we take ${\bm{F}}$ to be a smooth vector field on $\mathbb{R}^{d}$, and ${\bm{\Sigma}}(\bm{x})$ to be, for each $\bm{x}$, a $d\times d$ matrix with smooth entries, such that $(a_{ij}(\bm{x}))={\bm{\Sigma}}(\bm{x}){\bm{\Sigma}}^{T}(\bm{x})$ is a positive definite matrix for each $\bm{x}$ in $\mathbb{R}^{d}$.
Roughly speaking, a Lévy process $\bm{L}_{t}$ on $\mathbb{R}^{d}$ is a non-Gaussian stochastic process with independent and stationary increments that experience sudden, large jumps in any direction. The probability distribution of the these jumps is characterized by a non-negative Borel measure $\nu$ defined on $\mathbb{R}^{d}$ and concentrated on $\mathbb{R}^{d}\backslash\{0\}$ that satisfies the property $\int_{\mathbb{R}^{d}\backslash\{0\}}\min(1,\bm{y}^{2})\nu(\,\mathrm{d}\bm{y})<\infty$. This measure $\nu$ is called the jump measure of the Lévy process $\bm{L}_{t}$. Sometimes $X_{t}$ itself is referred to as a Lévy process with triplet $({\bm{F}},{\bm{\Sigma}},\nu)$. Within this convention, we reserve ourselves the terminology of a Lévy process to a process with triplet $(0,0,\nu)$. We refer to [149] and [148] for the mathematical background on Lévy processes.
Under suitable assumptions on ${\bm{F}}$, ${\bm{\Sigma}}$, and the Lévy measure $\nu$, the solution $X_{t}$ to Eq. (105) is a Markov process (e.g. [154]) and even a Feller process [155], with associated Kolmogorov operator taking the following integro-differential form for (e.g.) $\psi$ in $C^{\infty}_{c}(\mathbb{R}^{d})$ (Courrège theorem [156, 150]):
$\mathcal{L}_{K}\psi(\bm{x})$ $={\bm{F}}(\bm{x})\cdot\nabla\psi+\sum_{i,j=1}^{d}a_{ij}(\bm{x})\partial_{ij}\psi+\Gamma\psi(\bm{x}),\mbox{with}$ (106)
$\Gamma\psi(\bm{x})=\int_{\mathbb{R}^{n}\backslash\{0\}}\Big{[}\psi(\bm{x}+\bm{y})-\psi(\bm{x})-\bm{y}\cdot\nabla\psi(\bm{x})\mathds{1}_{\{\left\lVert\bm{y}\right\rVert<1\}}\Big{]}\nu(\,\mathrm{d}\bm{y}),$
where $\mathds{1}_{\{\left\lVert\bm{y}\right\rVert<1\}}$ denotes the indicator function of the (open) unit ball in $\mathbb{R}^{n}$; see also [149, Theorem 3.3.3].
The first-order term in Eq. (106) is the drift term caused by the deterministic, nonlinear dynamics. The second-order differential operator represents the diffusion part of the process $X_{t}$. It is responsible for the continuous component of the process.
The $\Gamma$-term, involving the integral, represents the jump part of the process. It captures the discontinuous jumps that the process experience due to the sudden changes caused by the Lévy process $\bm{L}_{t}$. Its intuitive interpretation breaks down as follows. The term, $\psi(\bm{x}+\bm{y})-\psi(\bm{x})$, calculates the difference in the test function value before and after the potential jump, capturing the change in the test function due to the jump.
The term $-\bm{y}\cdot\nabla\psi(\bm{x})\mathds{1}_{\{\left\lVert y\right\rVert<1\}}$ represents a first-order correction for small jumps. It aims to account for the fact that a small jump might not land exactly on the grid point ($\bm{x}+\bm{y}$), but somewhere in its vicinity. This term is often referred to as the Girsanov correction. Thus, the integral term $\Gamma$ accounts for all possible jump sizes ($\bm{y}$) within a unit ball, as weighted by the Lévy measure $\nu(\,\mathrm{d}\bm{y})$. The notion of Kolmogorov operators can be extended to SPDEs driven by cylindrical Wiener processes, with $\bm{x}$ lying in a Hilbert space; see [157]. Building up on such a functional framework, it is possible to provide a rigorous meaning of the Kolmogorov operator $\mathcal{L}_{K}$ in the case of an SPDE driven by a general Lévy process, forcing possibly infinitely many modes.
In the case of SPDE driven by a scalar jump process (forcing only two modes) studied in Section VI (see Eq. (94)), the jump measure is simple to describe. It corresponds to the measure $\nu(\,\mathrm{d}s)$ ($s$ real) associated with the real-valued jump process $f(t)$ given by Eq. (95). This measure is the Dirac delta $\lambda\delta_{s=1}$, where $\lambda$ is the intensity of the Poisson process associated with the random comb signal. Recall that this intensity is the limit of the probability of a single event occurring in a small interval divided by the length of that interval as the interval becomes infinitesimally small. Therefore, $\lambda$ is equal to the firing rate $f_{r}$ of $f(t)$.
In the case of SDEs driven by $\alpha$-stable Lévy processes,
$\nu(\,\mathrm{d}\bm{y})=c_{d,\alpha}\|\bm{y}\|^{-d-\alpha}\,\mathrm{d}\bm{y},$ (107)
where the non-Gaussianity index $\alpha$ lies in $(0,2)$, and $c_{d,\alpha}$ is a constant that depends on the dimension and involves the gamma function of Euler [152]. A $2$-stable ($\alpha=2$) process is simply a Brownian motion.
Figure 16: Chart presenting the main steps, objectives, and characteristics of the OPM reduction approach.
The rigorous theory of slow and center manifolds are in their infancy for SDEs driven by $\alpha$-stable Lévy processes [84, 158], but the main formulas share the same ingredients as recalled in Section II.2, with analogous backward-forward interpretations of those discussed in Section II.3 for the Lyapunov-Perron integrals involved therein ([84, Sec. 4]). The non-Markovian parameterizations of Section III and their data-driven optimization, can thus be extended, at least formally, for the case of SDEs driven by $\alpha$-stable Lévy processes and will be pursued elsewhere on concrete applications.
VIII A Chart Summarizing the Approach
To streamline the application of the OPM reduction approach for stochastically forced systems presented in this paper, we provide a concise overview of its key steps, objectives, and characteristics in the chart shown in Figure Fig. 16. This summary is tailored to the class of SPDEs discussed in Section II.1, with the primary goal of performing a reduction with a cutoff scale just below the forcing scale. This reduction involves a solution splitting where only the disregarded ”small” scales are stochastically forced (Figures 16A and 16B).
Given this splitting, where the disregarded scales are those directly forced by stochasticity, our approach reveals the importance of non-Markovian terms in the parameterization of these scales. These terms depend on the noise’s path history, enabling the tracking of small-scale fluctuations. The accuracy of this tracking is primarily governed by the memory content ($\tau$) of the parameterization, which is optimized from data by minimizing a natural loss function (Figure 16E) over a single noise realization $\omega$.
Once optimized, the resulting non-Markovian parameterization, the Optimal Parameterizing Manifold (OPM), allows us to capture the average behavior of small-scale fluctuations as a functional of large scales by averaging over noise paths (see inset below Figure 16B). To efficiently optimize the memory content ($\tau$) and avoid the computation of cumbersome quadratures, we decompose the $J$-term in Figure 16, $J(t,\bm{x})=\sum J_{n}(t){\bm{e}}_{n}(\bm{x})$, into random coefficients $J_{n}$ that solve simple auxiliary ODEs dependent on $\tau$ and the noise path $\omega$. The resulting OPM closure, exemplified by Equation (85), takes the form of an SDE system with coefficients that are nonlinear functionals of these $J_{n}$ terms (Figure 16F) reflecting the nonlinear interactions between the noise and nonlinear terms in the original system.
As demonstrated with SPDEs driven by Gaussian noise (Section V) or jump processes (Section VI), the OPM closure effectively predicts noise-induced transitions and rare event statistics. This success is attributed to our hybrid approach, which combines analytical insights for constructing the parameterizing manifold $\Phi$ (Figure 16C and Section III.1) with data-driven learning of the optimal memory content $\tau$ (Section III C).
IX Discussion and Perspectives
@@
Thus, this work demonstrates that Optimal Parameterizing Manifolds (OPMs) constructed from Backward-Forward (BF) systems and optimized using a single solution path can accurately reproduce noise-induced transitions across a broad range of out-of-sample, test paths. This approach is effective for non-Gaussian noise with jumps. A remarkable feature of the approach is that, even though these two different types of noise typically lead to different dynamical responses, the OPMs in these two settings take exactly the same structural form with the corresponding non-Markovian terms depending on the past history of the corresponding noise forcing the original SPDE; cf. Eqns. (40)–(41) and cf Eqns. (99)–(100).
@@
The formulas derived from the BF systems in Section III enable practical approximation of the theoretical optimal parameterization, particularly its non-Markovian random coefficients. These coefficients become crucial, especially when noise affects unresolved modes. They essentially encode how noise propagates to the resolved modes through nonlinear interactions.
@@
These more general stochastic parameterizations have the potential to significantly enhance the parameterization capabilities of $\Phi$ (Eq. (33)). This is because they encode higher-order interaction terms with respect to the noise amplitude, as shown in Eq. (30). Resolving these interactions is potentially important for improving the prediction of rare event statistics by non-Markovian reduced systems, especially when the noise intensity is low. This opens doors to potential applications using large deviation or rare event algorithms (e.g., [159, 160, 161, 162, 130, 163, 108, 21]). Additionally, iterating the BF system (Eq. (25)) beyond $\ell=2$ can also introduce higher-order terms involving the low-mode amplitude $X$.
The inclusion of higher-term in $X$ is non-unique though. For instance, inspired by [29, Sec. 7.2.1], stochastic parameterizations accounting for higher-order terms in $X$ are produced by solving analytically the following BF system:
@@
for which the two first equations for $p^{(1)}$ and $p^{(2)}$ are simultaneously integrated backward over $[t-\tau,t]$, followed by the forward integration of the $q_{n}$-equation over the same interval. Compared to Eq. (LABEL:Eq_BF_SPDEcubic), this BF system has two-layer in the low-mode variable $p$ resulting into parameterizations of order 4 in $X$; see [164, Theorem 2] for an example. These higher-order terms come from the nonlinear self-interactions between the low modes as brought by the term $\Pi_{c}G_{2}\big{(}p^{(1)}\big{)}$. Accounting for additional interactions either through (LABEL:Eq_h2) or the iterative procedure of Section II.3 can help improve the parameterization skills (see also [103, 134]), and should be considered if accuracy is key to resolve.
An alternative route to enhance accuracy would involve to parameterize stochastically the parameterization residual from the loss functions in (53), post-minimization. This technique has proven effective in modeling the fast motion transversal to OPM manifolds by means of networks of nonlinear stochastic oscillators [47, 43], as diagnosed by Ruelle-Pollicott resonances [165, 166]. Similarly, rapid dynamical fluctuations, transversal to (Markovian) OPM manifolds (Figure 9), could benefit from refined parameterizations through the incorporation of additional stochastic components following [47], potentially leading to more accurate representation of rare event statistics by the corresponding closures (Figure 7).
Extending beyond SPDEs driven by Markov processes, our framework can naturally accommodate SPDEs driven by non-Markovian stochastic processes. A prime example is when the stochastic forcing is a fractional Brownian motion (fBm) [167], for which the existence of unstable manifolds has been demonstrated [168]. By leveraging Lyapunov-Perron techniques, as employed in the construction of these manifolds (Section II.2), our extension program to parameterization can proceed analogously to Section III. This involves here as well, finite-time integration of relevant backward-forward systems to derive optimizable parameterizations from the governing equation. The random coefficients in these parameterizations would then become again functionals of the past history of the forcing noise, albeit this time incorporating a higher degree of memory compared to the stochastic process $Z^{n}_{\tau}$ used in the parameterizations Section III, due to the original memory content of the original forcing. Such generalized parameterizations should provide valuable insights into the interplay between nonlinearity and noise in physical systems exhibiting long-term memory effects [34, 35, 169].
Regardless of the Markovian or non-Markovian nature of the stochastic forcing, our reduction approach and its potential extensions offer valuable tools for addressing critical questions. By elucidating the intricate interplay between noise and nonlinearity, we can gain deeper dynamical insights. For example, we anticipate applying this reduction program to investigate early warning indicators of tipping phenomena, from a dynamical perspective, by tracing the energy flow from forced scales to physically relevant modes [170, 171]. Climate tipping elements provide a compelling domain for exploring these concepts and applying our methods [67, 172, 173].
The proposed stochastic parameterization method offers a new framework for understanding noise-nonlinearity interactions in stochastic multiscale systems. It is important to note that alternative techniques based on stochastic time scale separation, such as the stochastic path perturbation approach, have been employed to study related phenomena, including the random escape of stochastic fields from criticality in explosive systems [174] and first passage time problems in nonlocal Fisher equations [175, 176]. Future research directions could explore the potential of our method to obtain analytical approximations for first passage times in nonlocal SPDEs near criticality. Such an extension could provide valuable insights into pattern formations for a wide range of physical and biological problems such as in infectious disease modeling [177], nonlinear optics [178], or vegetation models [179].